# SAGA: Learning Signal-Aligned Distributions for Improved Text-to-Image Generation

### v1

@@ Abstract
State-of-the-art text-to-image models produce visually impressive results but often struggle with precise alignment to text prompts, leading to missing critical elements or unintended blending of distinct concepts. We propose a novel approach that learns a high-success-rate distribution conditioned on a target prompt, ensuring that generated images faithfully reflect the corresponding prompts. Our method explicitly models the signal component during the denoising process, offering fine-grained control that mitigates over-optimization and out-of-distribution artifacts. Moreover, our framework is training-free and seamlessly integrates with both existing diffusion and flow matching architectures. It also supports additional conditioning modalities -- such as bounding boxes -- for enhanced spatial alignment. Extensive experiments demonstrate that our approach outperforms current state-of-the-art methods. The code is available at https://github.com/grimalPaul/gsn-factory.
@@ 1 Introduction
We propose SAGA: Signal-Aligned Gaussian Approximation, a novel approach that generalizes the GSN framework to distribution learning within diffusion [21] and flow matching [29] models, enhancing text-to-image alignment. While GSN approaches shift an existing latent $\mathbf{z}_{t}$ towards a point better aligned with a prompt $\mathbf{y}$, our method learns a distribution $p(\mathbf{z}_{t}|\mathbf{y})\approx\mathcal{N}\left(\bm{\mu}_{\mathbf{y}},\bm{\Sigma}_{\mathbf{y}}\right)$ from which one can sample latents $\mathbf{z}_{t}$ that are well-aligned with the prompt $\mathbf{y}$. The mean vector $\bm{\mu}_{\mathbf{y}}$ explicitly captures the central signal component, providing superior control over the generative process and limiting saturated, out-of-distribution samples. Additionally, through extensive experimentation, we show that our method leads to improved sample generation with better adherence to the intended semantics, establishing its superiority over state-of-the-art approaches in generating semantically accurate images.
Another key advantage is that once this distribution is learned, we can efficiently sample multiple examples without requiring a separate optimization process for each $\mathbf{z}_{t}$, as it is necessary in GSN methods. This is particularly advantageous for real-world applications that often generate several synthetic images internally before selecting the best candidate via downstream selection mechanisms. While many state-of-the-art approaches rely on an external model to guide the sampling, our approach relies solely on the model’s internal knowledge, ensuring consistency in the denoising process while remaining training-free and seamlessly integrable with existing methods. Furthermore, it can incorporate other conditioning modalities, such as bounding boxes, to refine spatial alignment and improve generation quality further.
@@ 2 Related Work
Previous works [33, 41] demonstrated the coarse-to-fine nature of diffusion models, where low-frequency structures (the main scene) are reconstructed first, followed by finer details. Others [2, 33] further highlighted that textual guidance has the strongest influence at the early stages of generation, where the core elements of the scene are formed, making semantic information crucial. Lin et al. emphasized the phenomenon of signal leakage, which may explain why certain types of noise perform better than others [17]. This leakage can be further exploited [13] to control generated images’ style, brightness, and colors. More recently, the concept of winning tickets was introduced [23], which correspond to specific pixel blocks in the initial latent space that are particularly effective for generating certain concepts. By analyzing how the model perceives these pixel blocks and organizing them into bounding boxes per entity, they construct a latent representation that better aligns with the user’s intended scene, enabling finer control over object placement in the final image. InitNO [18] focuses on the issues of under-optimization and over-optimization in the GSN approach due to adjustments across multiple denoising steps. As shown in LABEL:fig:interp_latent_InitNO, they propose optimizing the initial noise at the first sampling step (at $t=T$) to ensure a more effective latent initialization, facilitating better image generation while preserving the original noise distribution. This approach samples from specific regions of the prior Gaussian distribution that minimize the risk of generating out-of-distribution samples. Our method enables learning a valid distribution that outperforms InitNO while effectively conditioning the latent representations based on a given prompt. Additionally, it supports conditioning with bounding boxes, similar to the lottery ticket hypothesis.
@@ 3 Preliminaries
These approaches construct a probabilistic path from $p_{\text{source}}$ to $p_{\text{target}}$. The forward process is defined by a conditional Gaussian distribution $p_{t}(\mathbf{z}_{t}\mid{}\mathbf{z}_{0})=\mathcal{N}(\mathbf{z}_{t};a_{t}\mathbf{z}_{0},b_{t}^{2}\mathbf{I})$, where $a_{t}$ and $b_{t}$ are schedule-dependent coefficients. This defines a trajectory of latent variables $\mathbf{z}_{t}$ for $t\in[0,T]$, where $t=0$ corresponds to the clean data latent $\mathbf{z}_{0}$ and $t=T$ to pure noise. The time $t$ is discrete for SD 1.4 and continuous for SD 3. Using the reparameterization trick, this path is written as $\mathbf{z}_{t}=a_{t}\mathbf{z}_{0}+b_{t}\bm{\epsilon}$, with $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. This process can be interpreted as interpolating the original signal $\mathbf{z}_{0}$ and noise $\bm{\epsilon}$, where $b_{t}$ gradually amplifies the noise part and obscures the signal, while $a_{t}$ attenuates the signal component.
@@ 4.1 Methodology and Formulation
Approximating $p(\mathbf{z}_{t}|\mathbf{y})$. Let us consider a visual generative model in which the forward process can be entirely defined by two sequences $(a_{t})_{t\in\{0,\dots,T\}}$ and $(b_{t})_{t\in\{0,\dots,T\}}$ of positive real numbers as explained in Section 3. First, we compute an expansion of $p(\mathbf{z}_{t}|\mathbf{y})$ with respect to $a_{t}$ and $b_{t}$, given by Proposition 1, which proof is in Appendix A.
@@
Consider a generative model that produces latent representations $\mathbf{z}_{0}$ conditioned on a prompt $\mathbf{y}$. In the forward process, the latents are noised according to $\mathbf{z}_{t}=a_{t}\mathbf{z}_{0}+b_{t}\bm{\epsilon},\ \bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ where $(a_{t})_{t\in\{0,\dots,T\}}$ and $(b_{t})_{t\in\{0,\dots,T\}}$ are sequences of positive real numbers. Let $\bm{\mu}_{\mathbf{y}}=\mathbb{E}[\mathbf{z}_{0}|\mathbf{y}]$ and $\bm{\Sigma}_{\mathbf{y}}=\mathrm{Var}(\mathbf{z}_{0}|\mathbf{y})$. If $\underset{t\rightarrow T}{\lim}(a_{t})=0$, $\underset{t\rightarrow T}{\lim}(b_{t})\neq 0$, and the cumulants of $\mathbf{z}_{0}|\mathbf{y}$ are finite, then:
@@
   $\mathbf{z}_{t}=a_{t}\bm{\tilde{\mu}}_{\mathbf{y}}+b_{t}\epsilon$
@@
Our method provides two sampling strategies to generate $N$ images for a prompt $\mathbf{y}$. The first involves denoising a single latent $\mathbf{z}_{T}$ to an intermediate step $t$. This single $\mathbf{z}_{t}$ initializes a distribution $q(\mathbf{z}_{t}|\mathbf{y})$ from which we sample and fully denoise $N$ latents to obtain the final images $\mathbf{z}_{0}$. The second strategy, which we adopt to maximize sample diversity, starts with $N$ independent latents $\mathbf{z}_{T}$. Each is denoised to step $t$, and each of the $N$ resulting latents $\mathbf{z}_{t}$ initializes its own distinct distribution. We then draw one latent from each of these $N$ distributions and complete their denoising. While we use the latter approach, both strategies yield similar quantitative scores, as detailed in Appendix G.
@@ 4.2 Cross-modal Attention Criterion
where $\mathcal{C}$ represents all possible pairs of subject tokens $(m,n)$, and $\mathbf{M}^{n}_{i,j}$ denotes the attention map value at position $(i,j)$ for subject token $n$. These criteria ensure a well-defined signal with sufficient attention for each subject token while minimizing overlap, guiding the optimization towards constructing an effective distribution given a prompt $\mathbf{y}$. The final criterion is computed as $\mathcal{L}=\frac{\mathcal{L}_{1}+\mathcal{L}_{2}}{2}$. While we adopt this formulation, other criteria could have been used as well. In Appendix E, we detail the extraction of attention features and the application of the criterion, along with additional processing steps for SD 1.4 and SD 3.
Figure 3 compares the attention maps from standard SD 1.4 with those from our method. The standard model activates only one entity’s map, failing to represent both concepts simultaneously. In contrast, our optimization procedure ensures that both attention maps are properly activated, demonstrating the successful capture of all entities specified in the prompt.
@@ 4.3 Rescaling the Signal
While our Gaussian formulation for $q(\mathbf{z}_{t}|\mathbf{y})$ is supported by a rigorous probabilistic framework, it does not guarantee that the final image $\mathbf{x}_{0}$ will be valid. In practice, the optimization can lead to over-saturated outputs, depending on the hyperparameter selection; the alignment score approximates the target condition but fails to enforce a natural dynamic range. As illustrated in Figure 4, an image generated without further control displays recognizable content yet suffers from excessive contrast and saturation. To control this, we adopt a rescaling strategy inspired by previous work [27]. After each optimization step, we rescale the learned signal $\bm{\tilde{\mu}}_{\mathbf{y}}$ to ensure its standard deviation does not exceed that of the $\mathbf{\hat{z}}_{0}$ used for initialization. Unlike the sample-wise guidance in GSN, this gives us direct control over the global properties of the learned signal. Further details and alternative methods are discussed in Appendix C.
@@ Methods.
Generation is done using 50 sampling steps for the SD 1.4 backbone and 28 sampling steps for SD 3. The distribution parameters are optimized using SGD with a learning rate of $20$ for both models. We apply 50 optimization steps for a fair comparison with GSN approaches. For tasks conditioned on bounding boxes, we replace the standard $\mathcal{L}_{2}$ loss with a dedicated criterion (detailed in Section E.1). This same criterion is then used both to learn the distribution in SAGA and for the subsequent GSN guidance in SAGA+. We selected all hyperparameters on a validation set. A detailed description of this set, a complete list of hyperparameters, and further ablations are provided in Appendix F. We present a key study on the sampling step $t$ in Section 5.4. On SD 1.4, we compare various training-free approaches. Note that GSN approaches are implemented only for SD 1.4. Specifically, on SD 1.4, we compare our method against Attend&Excite [5], SynGen [40], BoxDiff [53], InitNO [18]. InitNO+ refers to the variant of InitNO incorporating a GSN guidance. For layout-to-image generation, we use BoxDiff [53], UGD [3], and Lottery Tickets [23] as baselines. On SD 3, we compare our method against the baseline model. Comprehensive implementation details for all the mentioned methods are available in Appendix E.
@@ 5.2 Quantitative results
Table 3: User study results, reported as selection percentages. The $\emptyset$ symbol denotes trials where no method was chosen. Best results are in bold.
@@
We conducted a user study with 37 participants comparing SAGA+ with top methods on SD 1.4 (InitNO+, Syngen) and SAGA with SD 3. The study involved two distinct tasks: semantic matching, where participants could select multiple images matching a caption or none, and preference selection, where they could choose only their single preferred image. As reported in Table 3, results show a strong user preference for our method (protocol in Appendix H).
@@ 5.3 Qualitative Results
Figure 5 presents examples of image generation using the same seeds and prompts for different methods with GSN guidance. In the first column, only SAGA+ generates two semantically accurate images. In the last column, which features a complex prompt with three entities, only SAGA+ successfully generates the images with all the requested entities. Figure 6 presents the results of SAGA and SD 3. Our method improves text-to-image alignment, as suggested by the more consistent depiction of all entities in response to complex prompts. Further samples are provided in Appendix I.
@@ 5.4 Ablation Study
Figure 7: Performance of SAGA on SD 1.4 as a function of the sampling step. Each score is normalized. Points denote the mean performance averaged over various momentum values, with error bars indicating the standard deviation. The red curve highlights the Average Score.
@@ 7 Conclusion
In this work, we introduced a novel modeling approach to enhance image generation at inference time, offering new perspectives on controllable generation. By generalizing the GSN framework with a Gaussian-based formulation, we explicitly model the learned signal $\bm{\tilde{\mu}}_{\mathbf{y}}$, providing a new mechanism for guiding the generation process. Our experiments demonstrate the effectiveness of this approach in improving text-image alignment. Moreover, we show that our method applies to both flow matching and diffusion models. Future work could explore alternative forms of control over the learning of $\bm{\tilde{\mu}}_{\mathbf{y}}$.
@@ Appendix A Proof of 1
$=$ $\varphi(\mathbf{z}_{t};\bm{\mu},\bm{\Sigma})+O(a_{t}^{3})\quad$

### v2

@@ Abstract
State-of-the-art text-to-image models produce visually impressive results but often struggle with precise alignment to text prompts, leading to missing critical elements or unintended blending of distinct concepts. We propose a novel approach that learns a high-success-rate distribution conditioned on a target prompt, ensuring that generated images faithfully reflect the corresponding prompts. Our method explicitly models the signal component during the denoising process, offering fine-grained control that mitigates over-optimization and out-of-distribution artifacts. Moreover, our framework is training-free and seamlessly integrates with both existing diffusion and flow matching architectures. It also supports additional conditioning modalities -- such as bounding boxes -- for enhanced spatial alignment. Extensive experiments demonstrate that our approach outperforms current state-of-the-art methods. Code available at https://github.com/grimalPaul/gsn-factory
@@ 1 Introduction
We propose SAGA, Signal-Aligned Gaussian Approximation, a new approach that generalizes the GSN framework to distribution learning within diffusion [20] and flow matching [29] models, enhancing T2I alignment. While GSN approaches shift an existing latent $\mathbf{z}_{t}$ towards a point better aligned with a prompt $\mathbf{y}$, our method learns a distribution $p(\mathbf{z}_{t}|\mathbf{y})\approx\mathcal{N}\left(\bm{\mu}_{\mathbf{y}},\bm{\Sigma}_{\mathbf{y}}\right)$ from which one can sample latents $\mathbf{z}_{t}$ that are well-aligned with the prompt $\mathbf{y}$. The mean vector $\bm{\mu}_{\mathbf{y}}$ explicitly captures the central signal component, providing superior control over the generative process and limiting saturated, out-of-distribution samples. In addition, through extensive experimentation, we show that our method leads to improved sample generation with better adherence to the intended semantics, establishing its superiority over state-of-the-art approaches in generating semantically accurate images.
Another key advantage is that once this distribution is learned, we can efficiently sample multiple examples without requiring a separate optimization process for each $\mathbf{z}_{t}$, as is necessary in GSN methods. This is particularly advantageous for real-world applications that often generate several synthetic images internally before selecting the best candidate via downstream selection mechanisms. While many state-of-the-art approaches rely on an external model to guide the sampling, our approach relies solely on the model’s internal knowledge, ensuring consistency in the denoising process while remaining training-free and seamlessly integrable with existing methods. Furthermore, it can incorporate other conditioning modalities, such as bounding boxes, to refine spatial alignment and improve generation quality further.
@@ 2 Related Work
Previous works [33, 41] demonstrated the coarse-to-fine nature of diffusion models, where low-frequency structures (the main scene) are reconstructed first, followed by finer details. Others [2, 33] further highlighted that textual guidance has the strongest influence at the early stages of generation, where the core elements of the scene are formed, making semantic information crucial. Lin et al. emphasized the phenomenon of signal leakage, which may explain why certain types of noise perform better than others [17]. This leakage can be further exploited [13] to control generated images’ style, brightness, and colors. More recently, the concept of winning tickets was introduced [23], which corresponds to specific pixel blocks in the initial latent space that are particularly effective for generating certain concepts. By analyzing how the model perceives these pixel blocks and organizing them into bounding boxes per entity, they construct a latent representation that better aligns with the user’s intended scene, enabling finer control over object placement in the final image. InitNO [18] focuses on the issues of under-optimization and over-optimization in the GSN approach due to adjustments across multiple denoising steps. As shown in LABEL:fig:interp_latent_InitNO, they propose optimizing the initial noise at the first sampling step (at $t=T$) to ensure a more effective latent initialization, facilitating better image generation while preserving the original noise distribution. This approach samples from specific regions of the prior Gaussian distribution that minimize the risk of generating out-of-distribution samples. Our method enables learning a valid distribution that outperforms InitNO while effectively conditioning the latent representations based on a given prompt. Additionally, it supports conditioning with bounding boxes, similar to the lottery ticket hypothesis.
@@ 3 Preliminaries
These approaches construct a probabilistic path from $p_{\text{source}}$ to $p_{\text{target}}$. The forward process is defined by a conditional Gaussian distribution $p_{t}(\mathbf{z}_{t}\mid{}\mathbf{z}_{0})=\mathcal{N}(\mathbf{z}_{t};a_{t}\mathbf{z}_{0},b_{t}^{2}\mathbf{I})$ governed by the noise schedule of continuous, positive functions $a_{t},b_{t}$ on $t\in[0,T]$. This defines a trajectory of latent variables $\mathbf{z}_{t}$ for $t\in[0,T]$, where $t=0$ corresponds to the clean data latent $\mathbf{z}_{0}$ and $t=T$ to pure noise. Using the reparameterization trick, this path is written as $\mathbf{z}_{t}=a_{t}\mathbf{z}_{0}+b_{t}\bm{\epsilon}$, with $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. This process can be interpreted as interpolating the original signal $\mathbf{z}_{0}$ and noise $\bm{\epsilon}$, where $b_{t}$ gradually amplifies the noise part and obscures the signal, while $a_{t}$ attenuates the signal component.
@@ 4.1 Methodology and Formulation
Approximating $p(\mathbf{z}_{t}|\mathbf{y})$. Let us consider a visual generative model in which the forward process can be completely defined by $a_{t}$ and $b_{t}$, as explained in Section 3. First, we compute an expansion of $p(\mathbf{z}_{t}|\mathbf{y})$ with respect to $a_{t}$ and $b_{t}$, given by Proposition 1 (proof in Appendix A).
@@
Consider a generative model that produces latent representations $\mathbf{z}_{0}$ conditioned on a prompt $\mathbf{y}$. In the forward process, the latents are noised according to $\mathbf{z}_{t}=a_{t}\mathbf{z}_{0}+b_{t}\bm{\epsilon},\ \bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ where $a_{t}$ and $b_{t}$ are continuous, positive-valued functions defined on $t\in[0,T]$. Let $\bm{\mu}_{\mathbf{y}}=\mathbb{E}[\mathbf{z}_{0}|\mathbf{y}]$ and $\bm{\Sigma}_{\mathbf{y}}=\mathrm{Var}(\mathbf{z}_{0}|\mathbf{y})$. If $\underset{t\rightarrow T}{\lim}(a_{t})=0$, $\underset{t\rightarrow T}{\lim}(b_{t})\neq 0$, and the cumulants of $\mathbf{z}_{0}|\mathbf{y}$ are finite, then:
@@
   $\mathbf{z}_{t}=a_{t}\bm{\tilde{\mu}}_{\mathbf{y}}+b_{t}\bm{\epsilon}$
@@
Our method provides two sampling strategies to generate $N$ images for a prompt $\mathbf{y}$. The first involves denoising a single latent $\mathbf{z}_{T}$ to an intermediate step $t$. This single $\mathbf{z}_{t}$ initializes a distribution $q(\mathbf{z}_{t}|\mathbf{y})$ from which we sample and fully denoise $N$ latents to obtain the final images $\mathbf{z}_{0}$. The second strategy, which we adopt to maximize sample diversity, starts with $N$ independent latents $\mathbf{z}_{T}$. Each is denoised to step $t$, and each of the $N$ resulting latents $\mathbf{z}_{t}$ initializes its own distinct distribution. We then draw one latent from each of these $N$ distributions and complete their denoising. While we use the latter approach, both strategies yield similar quantitative scores (see Appendix G).
@@ 4.2 Cross-modal Attention Criterion
where $\mathcal{C}$ represents all possible pairs of subject tokens $(m,n)$, and $\mathbf{M}^{n}_{i,j}$ denotes the attention map value at position $(i,j)$ for subject token $n$. These criteria ensure a well-defined signal with sufficient attention for each subject token while minimizing overlap, guiding the optimization towards constructing an effective distribution given a prompt $\mathbf{y}$. The final criterion is computed as $\mathcal{L}=\frac{\mathcal{L}_{1}+\mathcal{L}_{2}}{2}$. While we adopt this formulation, other criteria could have been used as well. In Appendix E, we detail the extraction of attention features and the application of the criterion, along with additional processing steps for SD 1.4 and SD 3. Figure 3 compares the attention maps from standard SD 1.4 with those from our method. The standard model activates only one entity’s map, failing to represent both concepts simultaneously. In contrast, our optimization procedure ensures that both attention maps are properly activated, demonstrating the successful capture of all entities specified in the prompt.
@@ 4.3 Rescaling the Signal
While our Gaussian formulation for $q(\mathbf{z}_{t}|\mathbf{y})$ is supported by a rigorous probabilistic framework, it does not guarantee that the final image $\mathbf{x}_{0}$ will be valid. In practice, the optimization can lead to over-saturated outputs, depending on the hyperparameter selection; the alignment score approximates the target condition but fails to enforce a natural dynamic range. As illustrated in Figure 4, an image generated without further control displays recognizable content yet suffers from excessive contrast and saturation. To control this, we adopt a rescaling strategy inspired by [27]. After each optimization step, we rescale the learned signal $\bm{\tilde{\mu}}_{\mathbf{y}}$ to ensure its standard deviation does not exceed that of the $\mathbf{\hat{z}}_{0}$ used for initialization. Unlike the sample-wise guidance in GSN, this gives us direct control over the global properties of the learned signal. Further details and alternative methods are discussed in Appendix C.
@@ Methods.
We use 50 sampling steps for the SD 1.4 backbone and 28 sampling steps for SD 3. The distribution parameters are optimized using SGD with a learning rate of $20$ for both models. We apply 50 optimization steps for a fair comparison with GSN approaches. For tasks conditioned on bounding boxes, we replace the standard $\mathcal{L}_{2}$ loss with a dedicated criterion (detailed in Section E.1). This same criterion is then used both to learn the distribution in SAGA and for the subsequent GSN guidance in SAGA+. We selected all hyperparameters on a validation set. Details of this set and hyperparameters, as well as further ablations, are provided in Appendix F. We present a key study on the sampling step $t$ in Section 5.4. On SD 1.4, we compare various training-free approaches. With SD 1.4, we compare our method to GSN approaches Attend&Excite [5], SynGen [40], BoxDiff [53], InitNO [18]. InitNO+ refers to the variant of InitNO incorporating a GSN guidance. For layout-to-image generation, we use BoxDiff [53], UGD [3], and Lottery Tickets [23] as baselines. On SD 3, we compare our method to the baseline model. Comprehensive implementation details are available in Appendix E
@@ 5.2 Quantitative results
Table 3: User study results, reported as selection percentages. The $\emptyset$ symbol denotes trials where no method was chosen.
We conducted a user study with 37 participants comparing SAGA+ with top methods on SD 1.4 (InitNO+, Syngen) and SAGA with SD 3 (protocol in Appendix H). The study involved two distinct tasks: semantic matching, where participants could select multiple images matching a caption or none, and preference selection, where they could choose only their single preferred image. Users significantly prefer our method (Table 3).
@@ 5.3 Qualitative Results
We used the same seeds and prompts for different methods with GSN guidance to generate images. In Figure 5, the first column, only SAGA+ generates two semantically accurate images. In the last column, which features a complex prompt with three entities, only SAGA+ successfully generates the images with all the requested entities. Figure 6 presents the results of SAGA and SD 3. Our method improves text-to-image alignment, as suggested by the more consistent depiction of all entities in response to complex prompts. Further samples are provided in Appendix I.
@@ 5.4 Ablation Study
Figure 7: Performance of SAGA on SD 1.4 as a function of the sampling step. Each score is normalized. Points denote the performance averaged over various momentum values, with error bars indicating the standard deviation. The red curve highlights the Average Score.
@@ 7 Conclusion
We introduced a novel modeling approach to enhance image generation at inference time, offering new perspectives on controllable generation. By generalizing the GSN framework with a Gaussian-based formulation, we explicitly model the learned signal $\bm{\tilde{\mu}}_{\mathbf{y}}$, providing a new mechanism for guiding the generation process. Experiments showed that the approach improves text-image alignment for both flow matching and diffusion models. Future work could explore alternative forms of control over the learning of $\bm{\tilde{\mu}}_{\mathbf{y}}$.
Acknowledgments
This work was granted access to the HPC resources of IDRIS under the allocation 2022-AD011014009 made by GENCI. It relied on the use of the FactoryIA supercomputer, financially supported by the Ile-de-France Regional Council and was partly supported by the ANR-21-CE23-0024 IDeGeN project and the ANR-23-PEIA-0008 SHARP project in the context of the France 2030 program.
@@ Appendix A Proof of 1
$=$ $\varphi(\mathbf{z}_{t};\bm{\mu},\bm{\Sigma})+O(a_{t}^{3})\qquad$